OpenStudy (swissgirl):

\(\large \color{ Blue }{\frac{d^2y}{dx^2}+5 \left( \frac{dy}{dx}\right)^3-4y=e^x\)

OpenStudy (swissgirl):

@saifoo.khan

OpenStudy (saifoo.khan):

\[\large \color{blue}{\frac{d^2y}{dx^2}+5 \left( \frac{dy}{dx}\right)^3-4y=e^x}\]

OpenStudy (saifoo.khan):

You forgot to add '}' in the most end.

OpenStudy (swissgirl):

\(\large \color{ Blue }{\frac{d^2y}{dx^2}+5 \left( \frac{dy}{dx}\right)^3-4y=e^x}\) hmmmm it works here but not on my program :(

OpenStudy (swissgirl):

I just reset the page and it worked now. Its some stupid app that misbehaves

OpenStudy (saifoo.khan):

\large \color{blue}{\frac{d^2y}{dx^2}+5 \left( \frac{dy}{dx}\right)^3-4y=e^x}

OpenStudy (saifoo.khan):

Oh. Maybe.